OpenStudy (anonymous):

Louis is tilling the rectangular fool measuring 8.6 ft by 12 ft. how many 3by3 inch tile will Louis need to tile the floor with out over lapping? Guys i know the answer(408) i just need to figure out how to get it thank you!

Directrix (directrix):

A = l * w A = 8.6 (12) A = 103.2 square feet A 3 inch by 3 inch square tile is 3/12 feet by 3/12 feet. Or, 1/4 foot by 1/4 foot = 1/16 square foot of area which is .0625. 103.2 square feet divided by .0625 square feet = 1651.2 of the small tiles are required